ear 2 income statement data for Blue Hamster, then answer the questions that follow. Be sure to round each dollar value to the neares: whole dollar. Blue Ha Manufacturing Inc. Income Statement for Year Ending December 31 Year 2 (Forecasted) Year 1 Net sales $15,000,000 11,250,000 600,000 $3,150,000 315,000 2,835,000 1,134,000 $1,701,000 150,000 1,551,000 425,250 $1,125,750 Less: Operating costs, except depreciation and amortization Less: Depreciation and amortization expenses Less: Interest expense Less: Taxes (40%) Less: Preferred stock dividends Less: Common stock dividends 600,000 Operating income (or EBIT) Pre-tax income (or EBT) Earnings after taxes Earnings available to common shareholders Contribution to retained earnings $1,413,469
150,000 1,551,000 425,250 $1,125,750 Earnings available to common shareholders Less: Common stock dividends Contribution to retained earnings $1,413,469 Given the In Year 2, if Blue Hamster has 10,000 shares of preferred stock issued and outstanding, then each preferred share shoud in annual dividends. If Blue Hamster has 200,000 shares of common from stock issued and outstanding, then the firms earnings per share (EPS) is expected to change · in Year 2. in Year 1 to Blue Hamsters before interest, taxes, depreciation and amortization (EBITDA) value changed from in Year 1 to in Year 2 . It is to say that Blue Hamsters net inflows and outflows of cash at the end of Years 1 and 2 are equal to the companys annual of the item reported in the contribution to retained earnings, $1,125,750 and $1,413,469, respectively. This is because income statement involve payments and receipts of cash.
Answer #1
Fuzzy Button Clothing Company
Income statement for Year ending December 31
Year 1 Year 2 (Forecasted) Let Year2 sales = x, then
Net Sales 15000000 16918128 x-0.75x-600000=3629532
Less: Operating costs (75% of sales) 11250000 12688596 0.25x=4229532 => x=16918128
Less: Dep and amort. Expenses 600000 600000
Operating Income/EBIT 3150000 3629532
Less: Interest expense 315000 315000
Pre-tax income/EBT 2835000 3314532
Less: Taxes 40% 1134000 1325813
Earnings After taxes 1701000 1988719
Less: Preferred Stock Dividends 150000 150000
Earnings available to common shareholders 1551000 1838719
Less:Common Stock dividends 425250 425250
Contribution to retained earnings 1125750 1413469
1) Preferred shares outstanding 10000, dividend per PS = 150000/10000 = $15/ps
2)Common Shares outstanding 200000, change in EPS, Year 1 = 1551000/200000 = $7.76 to Year 2 = 1838719/200000 = $9.19.
3) EBITDA in Year 1 = 3150000 + 600000 = 3750000, in Year 2 = 3629532 + 600000 = 4229532
4) No, It is NOT to say----,                     because income statement do not show cash flow, as non-cash item, depreciation and
amortisation is a part of Income statement.